OpenStudy (anonymous):

can someone guide me to do this quiz please 10t/[3t^2]^2

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

I can explain this, but some responses would be great.

OpenStudy (anonymous):

Start by squaring the bottom value.

OpenStudy (unklerhaukus):

\[(ab^n)^m=a^mb^{n\times m}\] \[\frac{a^p}{a^q}=a^{p-q}\]

OpenStudy (unklerhaukus):

\[a=a^1\]